### v1

# Phonon assisted light absorption and emission in cubic-Boron Nitride

Abstract
Cubic boron nitride (cBN) is a wide–bandgap polymorph of boron nitride whose optical response remains only partially understood due to the coexistence of indirect electronic transitions and strong exciton–phonon coupling. Using first-principles many-body perturbation theory, we investigate the optical properties of cBN by combining GW quasiparticle corrections with Bethe-Salpeter equation calculations of excitonic effects. Phonon-assisted absorption and emission processes are explicitly included through the exciton–phonon coupling formalism. We find that phonon–mediated optical transitions provide a dominant contribution to both absorption and luminescence spectra, partially reconciling the discrepancy between the theoretical optical gap ( eV) and experimental emission around 6–7 eV. Our results demonstrate the importance of including exciton–phonon interactions for the correct interpretation of experimental spectra, offering new insights into light emission in wide-bandgap materials.
I Introduction
Boron nitride (BN) exhibits a rich polymorphism, encompassing phases ranging from the common hexagonal (hBN) and cubic (cBN) structures to more complex polytypes.Cazorla and Gould (2019) A significant challenge in the study of these materials is the frequent coexistence of different phases and the difficulty in obtaining perfectly pure samples. This complexity underscores the need for precise characterization techniques. While Raman spectroscopy is a widely used characterization tool in materials science, its utility for BN polymorphs is often limited due to the material’s few active Raman modes and the deep UV band gap forcing nonresonant conditions. Consequently, luminescence spectroscopy emerges as a far more powerful and informative probe. The sensitivity of luminescence to electronic structure, defects, and stacking order makes it exceptionally well-suited for characterizing BN nanostructures, where it can reveal details that remain inaccessible to Raman scatteringJaffrennou et al. (2007). For example, luminescence has proven to be a versatile tool, successfully distinguishing between different stacking sequences in the hexagonal phaseZanfrognini et al. (2023) and probing the effects of strain.Lechifflart et al. (2022); Schué (2017)
The past few years have witnessed intense experimental and theoretical interest in the luminescence properties of BN polymorphs.Cassabois et al. (2021); Schué et al. (2019a); Elias et al. (2021a) A unifying feature of most BN phases is their indirect band gap, meaning that luminescence typically arises from phonon-assisted transitions. Remarkably, despite this indirect nature, the luminescence efficiency in hBN has been shown to rival that of direct band gap materials.Cassabois et al. (2016); Schué et al. (2019a) This high efficiency should be a consequence of the strong exciton–phonon coupling inherent to BN systems.Schué et al. (2019b); Elias et al. (2021b) Beyond hBN, luminescence has also been investigated in the wurtzite phase (wBN), though a direct comparison between experiment and theory remains challenging due to difficulties in synthesizing pure wBN.Silvetti et al. (2023) Even in the direct band gap monolayer, luminescence studies are central in the debates over the origin of emission features, whether from direct or indirect transitions.Rousseau et al. (2024); Liu et al. (2025); Lechifflart et al. (2023); Marini et al. (2024); Fu et al. (2025)
Within this family of polymorphs, cubic boron nitride (cBN) presents a particularly compelling and unresolved puzzle.Samantaray and Singh (2005); Mirkarimi et al. (1997) As a thermodynamically stable phase at standard conditions, cBN has been synthesized primarily by chemical vapor deposition.Mirkarimi et al. (1997); Samantaray and Singh (2005) Experimental optical measurements, including those under pressure,Onodera et al. (1993) consistently report an optical absorption onset around 6–7 eV.Miyata et al. (1989); Galambosi et al. (2001); Sobolev and Zlobina (1999) However, state-of-the-art theoretical calculations, which typically neglect the electron–phonon interaction, predict a much larger direct optical gap of approximately 11 eV.Surh et al. (1991); Cappellini et al. (2001); Satta et al. (2004) This significant discrepancy has been tentatively attributed to phonon-assisted indirect transitions, which could lower the observed absorption edge. Furthermore, the potential presence of hBN inclusions in experimental samples, which luminesce around 6 eV, may mask the intrinsic phonon-assisted emission signal from pure cBN.
From a theoretical point of view, exciton–phonon coupling has been mainly studied with different approaches: by finite difference methodsCannuccia et al. (2019); Paleari et al. (2019); Silvetti et al. (2023); Lechifflart et al. (2022) and the direct calculation of microscopic exciton–phonon scattering amplitudes.Lechifflart et al. (2023); Marini et al. (2024); Zanfrognini et al. (2023) In this work, we follow this second strategy to investigate the role of atomic vibrations in the optical response of cBN.
This work aims to address the long-standing discrepancy between theoretical predictions and experimental optical measurements in cubic boron nitride. While previous studies have investigated the optical properties of cBN either neglecting phonon effects Satta et al. (2004) or treating phonon-assisted absorption without excitonic interactions Iqbal et al. (2024), a unified first-principles description combining excitons and phonon scattering has so far been missing.
Here, we present a first-principles investigation of the optical absorption and luminescence of cBN that explicitly includes exciton–phonon coupling within a many-body perturbation theory framework. By combining GW quasiparticle corrections, Bethe–Salpeter equation calculations, and phonon-assisted optical processes, we assess the role of indirect excitonic transitions in shaping the optical spectra of cBN and clarify their relevance for the interpretation of experimental observations.
This information is critical, given that cBN is a large-gap material where strongly bound excitons are expected and the coupling between excitons and phonons is known to be profound in other BN polymorphs.Schué et al. (2019b); Elias et al. (2021b) Furthermore, understanding exciton–phonon coupling in cBN has broader implications, particularly in light of recent studies on small polaron formation in doped cBN systemsShirodkar et al. (2023) and the reported observation of the electron–phonon Fano effect in cBN.Zhu et al. (2023).
The manuscript is organized as follows: in Sec. II we summarize all theoretical methods used in our calculations including the excitonic properties and the coupling with phonons; Sec. II.3 contains all computational parameters and code references used in this work; in Sec. III we present our results and compare them with the existing literature and experimental measurements; finally, in Sec. IV we draw conclusions and discuss future perspectives for this work.
II Theoretical methods
The ground-state properties of cubic boron nitride are calculated within density functional theory (DFT), as implemented in the Quantum ESPRESSO package Giannozzi et al. (2017). The experimental lattice parameter of 3.615 Å reported in Ref. Knittle et al., 1989 is used throughout. Vibrational properties and electron–phonon matrix elements are obtained using density functional perturbation theory (DFPT) Giannozzi et al. (2017). All parameters that enter the calculation are reported in Sec. II.3. Following our investigation of the ground-state properties, we shifted our focus to exploring the electronic structure and optical properties of BN. To study the electronic bands of BN, we diagonalized the Kohn-Sham (KS) Hamiltonian. However, as DFT is known to have limitations in accurately describing the electronic structure, including underestimating the electronic gapMartin et al. (2016), we used Many-Body Perturbation Theory implemented in the Yambo code Sangalli et al. (2019) to correct the KS eigenvalues. By doing so, we calculated the quasiparticle bandstructure as:
where are the corrected quasiparticle energies, is the renormalization factor, the exchange correlation KS potential and the self-energy is calculated in the GW approximation Martin et al. (2016), at frequency . The screened interaction is calculated using the Godby-Needs plasmon–pole model.Stankovski et al. (2011)
Starting from the quasiparticle bandstructure we proceeded to calculate the optical response function, including the electron–hole interaction. To achieve this, we solved the Bethe-Salpeter Equation (BSE) Strinati (1988), which includes both local–field effects and screened electron–hole attraction. The BSE can be reformulated as an eigenvalue problem, which is defined by a two-particle Hamiltonian Martin et al. (2016); Gatti and Sottile (2013) written on a basis of electron–hole transitions . In this work, we solved the BSE in the Tamm-Dancoff approximation, including therefore only resonant transitions, , where the momentum is in the first Brillouin zone ( and denoting valence and conduction bands, respectively). The matrix elements of the BSE are written in the following manner:
where is the energy associated with the above mentioned transition from valence to conduction, is the Coulomb potential derived from the variation of the Hartree term, and is the screened electron–hole interaction derived from the screened exchange.Strinati (1988) The energies that enter the BSE include the quasiparticle corrections calculated within the approximation. After diagonalization of Eq. 2 we obtain the eigenvectors and eigenvalues that are the eigenstates and the energies of the excitons (labeled by the index ), the stationary state of the BSE. The exciton energies and wave-functions at are used to build up the macroscopic dielectric function as:
where are the excitonic dipoles, are the dipole matrix elements between the Kohn-Sham states and is a small broadening term added to mimic the resolution of experimental spectra. To speed up convergence of the BSE with the number of k-points, we used the double-grid approach described in Ref. Alliati et al., 2022, see Sec II.3 for details.
II.1 Phonon-assisted light absorption
For the phonon-assisted absorption/emission, we follow the approach derived in Ref. Lechifflart et al., 2023 with two modifications. First, the phase of the exciton–phonon matrix elements has been fixed using the code and the approach described in Ref. Nalabothula, 2025. Second, the definition of phonon-assisted transition dipoles has been corrected to be gauge consistent. This corresponds to including also the off-diagonal matrix elements of the exciton-phonon self-energy of Ref. Lechifflart et al., 2023. Then, the first order correction to the dielectric constant Eq. 3 induced by the coupling with phonons can be written as:
here, are the excitonic dipoles and are the phonon-assisted coupling dipole moments. Eq. 4 describes the satellites generated by the emission or absorption of phonons by an exciton, and they appear as a peak at the energy of the finite-momentum excitons plus or minus one phonon energy. The ‘’ signs refer to phonon emission or absorption processes. The appearance of satellites also induces a renormalization of the purely exciton dielectric constant, a term that we do not consider here because the effect is small, and we investigate the two effects separately. For more details, see Ref. Lechifflart et al., 2023; Marini et al., 2024. The phonon-assisted dipole moments can be written explicitly as
where are the exciton–phonon coupling matrix elements at momentum Lechifflart et al. (2023); Marini et al. (2024)(since luminescence is always mediated by the recombination of an intermediate optical exciton, depends on a single momentum corresponding to the phonon momentum).
The small imaginary part in the phonon-assisted dipole moments derives from the exciton–phonon self-energy, see Ref. Lechifflart et al., 2023 . We set its value to 1 meV, which is large enough to remove unphysical divergences and small enough not to affect the final optical spectrum. In this work, we treat the bound exciton as a well-defined quasiparticle and therefore ignore possible overscreening effects.Paleari and Marini (2022); Stefanucci and Perfetto (2025) This approximation has been shown to be excellent in previous studies on BN systems.
II.2 Phonon assisted luminescence
To apply Eqs. (3) and (4) to the light emission process, we use a steady-state approximation to obtain the luminescence via the van Roosbroeck-Shockley (RS) relationVan Roosbroeck and Shockley (1954); Bebb and Williams (1972); Paleari et al. (2019):
In this equation, the oscillator strength of the peak is given by the exciton–phonon coupling matrix elements multiplied by the exciton dipoles . Here and are the energies of the optical and finite-momentum excitons, respectively, while are the phonon energies. The occupation function is
Here, is the temperature-dependent phonon Bose-Einstein occupation function. As can be seen, corresponds to processes of phonon “emission” () — the integer value representing spontaneous phonon emission also at — while corresponds to processes of phonon “absorption” (). Therefore, describes “light” emission by recombining excitons mediated by either “phonon” absorption or emission.
Luminescence is technically an out-of-equilibrium process, but we can assume that for a very low density of excitations and in steady-state conditions, the exciton population can be approximately described by an equilibrium distribution evaluated at an effective temperature . Here, we use the Boltzmann distribution. Experimentally, tends to coincide with the lattice temperature more or less above 100 K, while at very low temperature ( K), may vary between 10–50 K. The parameter is an effective excitonic temperature, used in the Boltzmann distribution to model excitonic occupations. Finally, in a real luminescence calculation, we replace the delta function in Eq. (6) by a Lorentzian, with a smearing set to 0.025 eV.
II.3 Computational details
Ground state properties of the Kohn-Sham bandstructure were obtained using the Quantum-Espresso (QE) codeGiannozzi et al. (2017) using DOJO norm-conserving pseudo-potentialsvan Setten et al. (2018) and PBE-sol exchange correlation functionalPerdew et al. (2008) and a cutoff for the wave-functions of 80 Ry. The phonons were also calculated with the QE code, using a k-grid and a q-grid.
GW and BSE calculations were performed using the Yambo codeSangalli et al. (2019). We used a cutoff of 6 Ha and 80 bands for the dielectric constant, and the same number of bands to expand the Green’s function. In the Bethe-Salpeter equation, we included 3 valence and 3 conduction bands. To speed up convergence with k-point, we used the double-grid method as implemented in Ref. Alliati et al., 2022 with the fine k-point grid of k-points. Exciton–phonon coupling matrix elements were calculated using YamboPyPaleari et al. (2025) and the LetzElPhC libraryLetzElPhC using the same band range of the BSE and the same q-grid.
The symmetry and selection rule analysisNalabothula (2025) was also performed with Yambopy via the spglibTogo et al. (2024) and MolSymGoodlett et al. (2024) packages.
In the calculation of phonon-assisted light absorption, exciton and phonon energies are interpolated on a large q-grid and used in the double-grid version of Eqs. (4) and (6) as described in Appendix A of Ref. Lechifflart et al., 2023. In particular, phonon energies were interpolated using force constants from QEGiannozzi et al. (2017) while excitons were interpolated on the same q-grid using a smooth Fourier interpolation.Pickett et al. (1988) For the calculation of the phonon-assisted dielectric constant we included the lowest 9 excitonic states and we include 20 scattering states in the sum of Eq. 3, that is enough to converge the phonon-assisted absorption/emission, in the luminescence spectrum we consider an excitonic and phononic temperature of 10 K, corresponding to the occupation of the deep minimum of the exciton dispersion.
III Results
In this section, we present the results for the various properties of BN.
III.1 Electronic and phononic bandstructures
In the top panel of Fig. 1, we report the bandstructure of cBN calculated at the KS and G0W0 level. We find the direct band gap at of 8.82 eV and 10.91 eV at DFT and G0W0 levels. The indirect gap between is 4.34 eV and 6.0 eV. These values have to be compared with results from S. Louie et al. Surh et al. (1991) who found 4.3 eV and 6.3 eV for the indirect and 8.6 eV and 11.4 eV for the direct one. G. Satta et al. found 11.4 eV for the direct band gap Satta et al. (2004) and experiments report and 14.5 eV Philipp and Taft (1962). In general we expect our GW correction to be smaller than the one obtained from the S. Louie et al. due to the different plasmon–pole model (for a discussion see Ref. Stankovski et al., 2011), while regarding the comparison with G. Satta et al. two factors can explain the difference: first, we used the experimental lattice constant while they optimized it in DFT, and second, the parameters of our calculations are much more converged. Regarding the experimental result, it must be considered with great uncertainty, given that it is a marginal result with respect to the main aim of that work.
In general, it is well known that in large band gap materials G0W0 tends to underestimate the gap correction and usually self-consistency improves the agreement with the experimentsArtús et al. (2021). However, in this work, we decided to stick with the G0WBSE approximation in order to be consistent with other works in the literature and on the assumption that any possible error in this approximation is of the same order in all BN polymorphs, so that we can compare our results with others on similar materials. We also neglect all energy renormalizations due to electron-phonon interactions, namely quasiparticle band-gap reduction and exciton binding energy increase, which partially compensate. In summary, our results should be interpreted with the understanding that certain physical effects, mostly pertaining to the absolute energy position of the optical spectra, are not taken into account, while they provide a good description of the optical response of cBN as far as coupling strengths and relative energy positions are concerned.
In the bottom panel of Fig. 1, we report the phonon bandstructure calculated with Density Functional Perturbation Theory (DFPT) as implemented in the QE code. With the same approach, we also calculated the deformation potential that enters into the definition of the electron–phonon and then the exciton–phonon matrix elements. We found an LO-TO splitting at of 29 meV and the overall phonon bands are in good agreement with experimental measurements Cai et al. (2007); Eremets et al. (1995) and previous calculationsOhba et al. (2001).
III.2 Optical properties and exciton dispersion
Starting from the quasiparticle bandstructure calculated in the previous section, we solved the Bethe-Salpeter equation (Eq. (2)). Then we calculated the optical absorption of cBN according to Eq. 3. In agreement with previous studies Satta et al. (2004), we found that the optical response of cBN requires a very dense grid to converge. For this reason, we repeated the calculation using the Haydock technique (an iterative solver) combined with a double-grid approachAlliati et al. (2022) to obtain smoother spectra. In Fig. 2 we report the BSE results compared with the independent particle approximation (IPA) starting from the KS bandstructure.
We found that the onset of absorption begins at around 10.5 eV, at much higher energies than those reported in experiments.
In Fig. 2 we have also indicated the first two absorption peaks with letters A and B, the first being a bound state with binding energy of approximately 0.3 eV and the second above the G0W0 gap.
Compared to the IPA, the combination of GW+BSE shifts the spectrum, and several bound excitonic states form.
However, the situation is quite different from that of hBN, where all the spectral weight is concentrated in a few excitonic statesPaleari et al. (2018).
In the cBN case, the spectrum is the sum of many excitonic states with close energies. This is clearly visible when the excitonic dispersion is plotted.
To do this, we solved BSE at finite momentum on a regular q-grid (see Sec. II.3).
In Fig. 3, we report the exciton dispersion along high-symmetry lines. The indirect nature of cBN is clearly visible: in particular, the point (corresponding to the optical limit) does not lie at the bottom of a valley, but instead is one of the highest energy points in the bands.
Here, we have a low-energy subset of 9 exciton states between 10.54 and 10.6 eV. The first absorption peak visible in Fig. 2, denoted as ‘A’, lies in this subset: it is three times degenerate and transforms as the representation of the point group of cBN (only states with this symmetry satisfy the optical selection rules and can be bright).
Furthermore, the LT splitting of this state - verified by solving the BSE at with and without the long-range part of the Coulomb interaction - amounts to 0.05 eV (the longitudinal state jumps from 10.55 to 10.6 eV).
There is also a dark triplet lying 15 meV below the A peak, while the remainder of the 9-state subset is composed by a doubly degenerate state () and a nondegenerate one ().
The minimum of the excitonic dispersion is located at the point at eV, almost 5 eV below the point, consistent with the deep indirect nature of the electronic gap.
Here, the minimum is formed by three states: first, a doubly degenerate state transforming as the representation of the little group of the point, followed 5 meV above by a nondegenerate state.
III.3 Phonon-assisted light absorption/emission
In Fig. 4 we report the phonon-assisted luminescence of cBN calculated using Eq. 6 at 10 K. At this temperature, the lowest-energy exciton () at is occupied, while the state above remains empty. Indeed, the luminescence originates from excitons at the point. We found that the signal is centered around 5.55–5.6 eV, well below that of hBN (from 5.77–5.9 eV). In the same figure, we report the decomposition of the luminescence signal into the contribution of the different phonon modes.
To interpret the luminescence spectrum, we refer to the phonons shown in the lower panel of Fig. 1, highlighted by a vertical red line in the same figure. These are the phonon modes that match the momentum of the lowest excitonic states in Fig. 3. In order to distinguish the different phonons modes we use the same nomenclature as for phonons, as it is clear from the continuous lines of the phonon dispersion which mode we are referring to. Combining all this information we found that phonon-assisted transitions are primarily mediated by transverse phonon branches with a small contribution of longitudinal branches at lower energies.
For the exciton, it is important to emphasize that the selection rules for the phonon-mediated recombination process (which requires a bright exciton as final state, see Eq. 6) allow for coupling with any phonon mode. Instead, in the case of the exciton, the coupling with the mode corresponding to the ‘LA’ branch is forbidden.
Compared to the more famous and better characterized luminescence spectrum of hBNCannuccia et al. (2019); Paleari et al. (2019); Zanfrognini et al. (2023), the spectrum of cBN presents several differences: firstly, it is located at a lower energy; secondly, longitudinal and transverse phonons at are not so separate as in hBN to generate a doublet in luminescence; thirdly, unlike hBN, the phonon modes that contribute most are from the ‘TO’ and ‘TA’ branches, while ‘LA’ and ‘LO’ are low or irrelevant. All these characteristics will make it possible to distinguish it in future experiments on clean samples of cBN.
Finally, in Fig. 5 we report the low energy tail of the dielectric function including the contribution from the exciton–phonon scattering obtained from Eq. 4 interpolated on a double-grid for exciton and phonon energiesLechifflart et al. (2023) at 10 K. We found that coupling with phonons red–shifts the onset from 10.3 to 9.8 eV. Looking at the excitonic dispersion, Fig. 3, one would expect even a lower energy contribution of the epsilon. However, transitions toward low excitons close to have a very weak intensity due to the large denominators that appear in Eq. 5. For this reason, the first visible contribution to the epsilon is only 1 eV below the direct transitions.
IV Conclusions
In this work, we have investigated the optical response of cubic boron nitride (cBN) using first-principles many-body perturbation theory, specifically accounting for the critical role of exciton–phonon coupling.
Our findings provide a better understanding of the long-standing discrepancy between theoretical predictions of a eV optical gap and experimental reports of absorption and emission features in the 6-7 eV range.
We have shown that while the direct optical gap remains high, the inclusion of exciton–phonon interactions effectively lowers the observable absorption onset.
However, intrinsic exciton–phonon coupling in the pristine system is not enough to reconcile theory with experimental observations. In fact, even if excitons at very low energy are present in cBN, their contribution to the dielectric constant is suppressed by the large denominators in the phonon-assisted dipoles. We expect that the low-energy contribution in the experimental measurement of the dielectric constant of cBN is due to the presence of defects, through direct transition to defect levels or to the activation of transition to finite momentum excitons due to the breaking of translation symmetryTararan et al. (2018).
Regarding the luminescence, our results show that cBN and hBN exhibit distinct spectral profiles. We found that the intrinsic phonon-assisted luminescence of pure cBN is centered around 5.57–5.60 eV, which is lower than that typically observed for hBN, from 5.77–5.9 eV. This energy shift suggests that experimental emission signals reported around 6 eV may be attributed to hBN inclusions within cBN samples rather than the intrinsic cBN or defect–related states. This underscores the sensitivity of luminescence spectroscopy as a tool for phase characterization in boron nitride polymorphs.
Acknowledgements.
C.A. and A.P. acknowledge B. Demoulin and A. Saul for the management of the computer cluster Claudia. F.P. acknowledges Muralidhar Nalabothula for insightful discussions on exciton symmetries and the evaluation of luminescence spectra. C.A. acknowledges funding from European Research Council MSCA-ITN TIMES under grant agreement 101118915. E.C. acknowledges ANR project COLIBRI No. ANR-22-CE30-0027. C.A., A.M. and A.P. acknowledge support from the Amidex foundation through the project INDIGENA. F.P. Acknowledges support from ICSC - Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing - funded by the European Union through the Italian Ministry of University and Research under PNRR M4C2I1.4 (Grant No. CN00000013). This work has also been supported by the European Union through the MaX ”MAterials design at the eXascale” Centre of Excellence (Grant agreement No. 101093374, co-funded by the EuroHPC JU).References
- Cazorla and Gould (2019) C. Cazorla and T. Gould, Science Advances 5, eaau5832 (2019), https://www.science.org/doi/pdf/10.1126/sciadv.aau5832 .
- Jaffrennou et al. (2007) P. Jaffrennou, J. Barjon, J.-S. Lauret, B. Attal-Trétout, F. Ducastelle, and A. Loiseau, Journal of Applied Physics 102 (2007).
- Zanfrognini et al. (2023) M. Zanfrognini, A. Plaud, I. Stenger, F. Fossard, L. Sponza, L. Schué, F. Paleari, E. Molinari, D. Varsano, L. Wirtz, F. m. c. Ducastelle, A. Loiseau, and J. Barjon, Phys. Rev. Lett. 131, 206902 (2023).
- Lechifflart et al. (2022) P. Lechifflart, F. Paleari, and C. Attaccalite, SciPost Phys. 12, 145 (2022).
- Schué (2017) L. Schué, “Propriétés optiques et structurales du nitrure de bore en hybridation sp2: des cristaux massifs aux feuillets atomiques,” https://tel.archives-ouvertes.fr/tel-01827237/ (2017), ph.D. thesis, Université Paris-Saclay.
- Cassabois et al. (2021) G. Cassabois, A. Rousseau, C. Elias, T. Pelini, P. Vuong, P. Valvin, and B. Gil, Comptes Rendus. Physique 22, 1 (2021).
- Schué et al. (2019a) L. Schué, L. Sponza, A. Plaud, H. Bensalah, K. Watanabe, T. Taniguchi, F. Ducastelle, A. Loiseau, and J. Barjon, Phys. Rev. Lett. 122, 067401 (2019a).
- Elias et al. (2021a) C. Elias, G. Fugallo, P. Valvin, C. L’Henoret, J. Li, J. Edgar, F. Sottile, M. Lazzeri, A. Ouerghi, B. Gil, et al., Physical Review Letters 127, 137401 (2021a).
- Cassabois et al. (2016) G. Cassabois, P. Valvin, and B. Gil, Nature Photonics 10, 262 (2016).
- Schué et al. (2019b) L. Schué, L. Sponza, A. Plaud, H. Bensalah, K. Watanabe, T. Taniguchi, F. m. c. Ducastelle, A. Loiseau, and J. Barjon, Phys. Rev. Lett. 122, 067401 (2019b).
- Elias et al. (2021b) C. Elias, G. Fugallo, P. Valvin, C. L’Henoret, J. Li, J. H. Edgar, F. Sottile, M. Lazzeri, A. Ouerghi, B. Gil, and G. Cassabois, Phys. Rev. Lett. 127, 137401 (2021b).
- Silvetti et al. (2023) M. Silvetti, C. Attaccalite, and E. Cannuccia, Phys. Rev. Mater. 7, 055201 (2023).
- Rousseau et al. (2024) A. Rousseau, J. Plo, P. Valvin, T. S. Cheng, J. Bradford, T. James, J. Wrigley, C. Mellor, P. Beton, S. Novikov, et al., 2D Materials 11, 025026 (2024).
- Liu et al. (2025) L. Y. Liu, S. Y. Woo, J. Wu, B. Hou, C. Su, and D. Y. Qiu, arXiv preprint arXiv:2502.20454 (2025).
- Lechifflart et al. (2023) P. Lechifflart, F. Paleari, D. Sangalli, and C. Attaccalite, Phys. Rev. Mater. 7, 024006 (2023).
- Marini et al. (2024) G. Marini, M. Calandra, and P. Cudazzo, Nano Letters 24, 6017 (2024).
- Fu et al. (2025) L. Fu, Y. Hu, N. Tang, J. Duan, X. Jia, H. Yang, Z. Li, X. Han, G. Li, J. Lu, L. Dai, W. Ge, Y. Yao, and B. Shen, Phys. Rev. Lett. 135, 046903 (2025).
- Samantaray and Singh (2005) C. Samantaray and R. Singh, International Materials Reviews 50, 313 (2005).
- Mirkarimi et al. (1997) P. Mirkarimi, K. McCarty, and D. Medlin, Materials Science and Engineering: R: Reports 21, 47 (1997).
- Onodera et al. (1993) A. Onodera, M. Nakatani, M. Kobayashi, Y. Nisida, and O. Mishima, Physical Review B 48, 2777 (1993).
- Miyata et al. (1989) N. Miyata, K. Moriki, O. Mishima, M. Fujisawa, and T. Hattori, Physical Review B 40, 12028 (1989).
- Galambosi et al. (2001) S. Galambosi, J. Soininen, K. Hämäläinen, E. L. Shirley, and C.-C. Kao, Physical Review B 64, 024102 (2001).
- Sobolev and Zlobina (1999) V. Sobolev and M. Zlobina, Journal of Applied Spectroscopy 66, 654 (1999).
- Surh et al. (1991) M. P. Surh, S. G. Louie, and M. L. Cohen, Phys. Rev. B 43, 9126 (1991).
- Cappellini et al. (2001) G. Cappellini, G. Satta, M. Palummo, and G. Onida, Physical Review B 64, 035104 (2001).
- Satta et al. (2004) G. Satta, G. Cappellini, V. Olevano, and L. Reining, Physical Review B—Condensed Matter and Materials Physics 70, 195212 (2004).
- Cannuccia et al. (2019) E. Cannuccia, B. Monserrat, and C. Attaccalite, Phys. Rev. B 99, 081109 (2019).
- Paleari et al. (2019) F. Paleari, H. P.C.Miranda, A. Molina-Sánchez, and L. Wirtz, Phys. Rev. Lett. 122, 187401 (2019).
- Iqbal et al. (2024) S. Iqbal, T. Cheng, X. Duan, L. Liu, and J.-Y. Yang, Journal of Applied Physics 135 (2024).
- Shirodkar et al. (2023) S. N. Shirodkar, C. A. Sayou Ngomsi, and P. Dev, ACS Applied Electronic Materials 5, 1707 (2023).
- Zhu et al. (2023) S. Zhu, L. Hu, Z. Lin, X. Wei, J. He, and W. Zheng, ACS Photonics 11, 286 (2023).
- Giannozzi et al. (2017) P. Giannozzi, O. Andreussi, T. Brumme, O. Bunau, M. B. Nardelli, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, M. Cococcioni, N. Colonna, I. Carnimeo, A. D. Corso, S. de Gironcoli, P. Delugas, R. A. DiStasio, A. Ferretti, A. Floris, G. Fratesi, G. Fugallo, R. Gebauer, U. Gerstmann, F. Giustino, T. Gorni, J. Jia, M. Kawamura, H.-Y. Ko, A. Kokalj, E. Küçkbenli, M. Lazzeri, M. Marsili, N. Marzari, F. Mauri, N. L. Nguyen, H.-V. Nguyen, A. O. de-la Roza, L. Paulatto, S. Poncé, D. Rocca, R. Sabatini, B. Santra, M. Schlipf, A. P. Seitsonen, A. Smogunov, I. Timrov, T. Thonhauser, P. Umari, N. Vast, X. Wu, and S. Baroni, Journal of Physics: Condensed Matter 29, 465901 (2017).
- Knittle et al. (1989) E. Knittle, R. M. Wentzcovitch, R. Jeanloz, and M. L. Cohen, Nature 337, 349 (1989).
- Martin et al. (2016) R. M. Martin, L. Reining, and D. M. Ceperley, Interacting electrons (Cambridge University Press, 2016).
- Sangalli et al. (2019) D. Sangalli, A. Ferretti, H. Miranda, C. Attaccalite, I. Marri, E. Cannuccia, P. Melo, M. Marsili, F. Paleari, A. Marrazzo, et al., Journal of Physics: Condensed Matter 31, 325902 (2019).
- Stankovski et al. (2011) M. Stankovski, G. Antonius, D. Waroquiers, A. Miglio, H. Dixit, K. Sankaran, M. Giantomassi, X. Gonze, M. Côté, and G.-M. Rignanese, Phys. Rev. B 84, 241201 (2011).
- Strinati (1988) G. Strinati, Riv. Nuovo Cimento 11, 1 (1988).
- Gatti and Sottile (2013) M. Gatti and F. Sottile, Phys. Rev. B 88, 155113 (2013).
- Alliati et al. (2022) I. M. Alliati, D. Sangalli, and M. Grüning, Frontiers in Chemistry 9, 763946 (2022).
- Nalabothula (2025) M. Nalabothula, Symmetries of Excitons: Implications for Exciton-Phonon Coupling and Optical Spectroscopy, Ph.D. thesis, Unilu - Université du Luxembourg [The Faculty of Science, Technology and Medicine], Luxembourg, Luxembourg (12 May 2025), https://orbilu.uni.lu/10993/65236 .
- Paleari and Marini (2022) F. Paleari and A. Marini, Phys. Rev. B 106, 125403 (2022).
- Stefanucci and Perfetto (2025) G. Stefanucci and E. Perfetto, SciPost Phys. 18, 009 (2025).
- Van Roosbroeck and Shockley (1954) W. Van Roosbroeck and W. Shockley, Physical Review 94, 1558 (1954).
- Bebb and Williams (1972) H. B. Bebb and E. W. Williams, Semiconductors and semimetals (Elsevier, 1972) Chap. Photoluminescence I: Theory.
- van Setten et al. (2018) M. van Setten, M. Giantomassi, E. Bousquet, M. Verstraete, D. Hamann, X. Gonze, and G.-M. Rignanese, Computer Physics Communications 226, 39 (2018), http://www.pseudo-dojo.org/.
- Perdew et al. (2008) J. P. Perdew, A. Ruzsinszky, G. I. Csonka, O. A. Vydrov, G. E. Scuseria, L. A. Constantin, X. Zhou, and K. Burke, Physical review letters 100, 136406 (2008).
- Paleari et al. (2025) F. Paleari, A. Molina-Sánchez, M. Nalabothula, R. Reho, M. Bonacci, J. M. Castelo, J. Cervantes-Villanueva, M. Pionteck, M. Silvetti, C. Attaccalite, and H. Pereira Coutada Miranda, “YamboPy,” (2025).
- (48) LetzElPhC, “Lëtzebuerg Electron Phonon Code,” https://gitlab.com/lumen-code/LetzElPhC.
- Togo et al. (2024) A. Togo, K. Shinohara, and I. Tanaka, Science and Technology of Advanced Materials: Methods 4, 2384822 (2024), https://doi.org/10.1080/27660400.2024.2384822 .
- Goodlett et al. (2024) S. M. Goodlett, N. L. Kitzmiller, J. M. Turney, and I. Schaefer, Henry F., The Journal of Chemical Physics 161, 024107 (2024), https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/5.0216738/20039263/024107_1_5.0216738.pdf .
- Pickett et al. (1988) W. E. Pickett, H. Krakauer, and P. B. Allen, Phys. Rev. B 38, 2721 (1988).
- Philipp and Taft (1962) H. R. Philipp and E. A. Taft, Phys. Rev. 127, 159 (1962).
- Artús et al. (2021) L. Artús, M. Feneberg, C. Attaccalite, J. H. Edgar, J. Li, R. Goldhahn, and R. Cuscó, Advanced Photonics Research 2, 2000101 (2021).
- Cai et al. (2007) Y. Cai, L. Zhang, Q. Zeng, L. Cheng, and Y. Xu, Solid state communications 141, 262 (2007).
- Eremets et al. (1995) M. Eremets, M. Gauthier, A. Polian, J. Chervin, J. Besson, G. Dubitskii, and Y. Y. Semenova, Physical Review B 52, 8854 (1995).
- Ohba et al. (2001) N. Ohba, K. Miwa, N. Nagasako, and A. Fukumoto, Phys. Rev. B 63, 115207 (2001).
- Pouch and Alterovitz (1990) J. J. Pouch and S. A. Alterovitz, in Materials Science Forum, Vol. 54 (Trans Tech Publications Ltd, 1990) p. 165.
- Paleari et al. (2018) F. Paleari, T. Galvani, H. Amara, F. Ducastelle, A. Molina-Sánchez, and L. Wirtz, 2D Materials 5, 045017 (2018).
- Tararan et al. (2018) A. Tararan, S. Di Sabatino, M. Gatti, T. Taniguchi, K. Watanabe, L. Reining, L. Tizei, M. Kociak, and A. Zobelli, Physical Review B 98, 094106 (2018).